OpenStudy (anonymous):

solve for all values of x: log4(x^2 + 3x) - log4 (x+5)=1

OpenStudy (raden):

hint : log4(x^2 + 3x) - log4 (x+5) = log4 (x^2 + 3x)/(x+5) and 1 = log4 (4)

OpenStudy (anonymous):

thankss!!

OpenStudy (solomonzelman):

I love askers that only require a hint, as opposed to the answer!

OpenStudy (anonymous):

k